# Beyond Chain-of-Thought: A Survey of Chain-of-X Paradigms for LLMs

### v1

@@ 3.2 Chain-of-Augmentation
A popular variant of CoX methods is Chain-of-Augmentation, where the chain is augmented with additional knowledge. Based on the types of augmented data, we categorize them as follows.
@@
Instructions serve as an important augmentation, guiding LLMs through complex reasoning or task execution processes where determining the next step can be nontrivial Zha et al. (2023). For instance, the Chain-of-InstructEditing framework Zhang et al. (2023d) harnesses this concept by generating sequential instructions to guide image editing tasks, illustrating how specific editions can refine the output by focusing on relevant areas. Moreover, Zha et al. (2023) introduce the Chain-of-Command to tackle ambiguities in user instructions for table manipulations. Inferring from user instructions, it enables LLMs to employ a series of precise pre-defined commands for more accurate table execution. In the realm of e-commerce, Li et al. (2024b) implement a similar structured approach with their Chain-of-Task, which breaks down customer interactions into manageable atomic tasks, significantly simplifying complex operations. Similarly, the Chain-of-Instructions framework proposed by Hayati et al. (2024) iteratively solves decomposed subtasks using outputs of previous steps as instructions for the next step. The results show that stepwise guidance can substantially improve both the process and the outcomes of complex problem-solving tasks.
@@
Utilizing historical data for informed predictive modeling is another facet of Chain-of-Augmentation, drawing contextual insights from past interactions or events. This approach is exemplified by Do et al. (2023)’s Chain-of-Opinion, which analyzes historical user opinions to predict future reactions, offering valuable foresight into user sentiment. In user-interface exploration, Zhan and Zhang (2023) apply a Chain-of-Actiona framework, leveraging past actions to guide future interactions, thereby optimizing user experience through learned behaviors. Ma et al. (2023) take a similar approach in gaming environments like StarCraft II, where a Chain-of-Summarization provides strategic recommendations based on a synthesis of past gameplay observations. The development of taxonomy structures also benefits from historical data, as seen in the Chain-of-Layer by Zeng et al. (2024), which builds upon previously identified categories to enhance classification tasks. Temporal knowledge graphs receive a forward-looking treatment as well with methods like Chain-of-History from Luo et al. (2024) and Xia et al. (2024d), where historical graph structures inform predictions about future linkages and interactions.
Retrieval
Chain-of-Retrievals methods are designed to intersperse the generation process with sequences of explicit retrievals, enhancing the quality of the generated content Zhao et al. (2023). For instance, Xu et al. (2023) introduce the Chain-of-Query framework, which improves the search capabilities of LLMs through a systematic arrangement of query-answer pairs, each aimed at enhancing information retrieval. Similarly, Chain-of-Question proposed by Huang et al. (2024) focuses on refining query mechanisms, where each sub-question decomposed from the original question helps in retrieving more accurate knowledge from external knowledge base. Further refining this concept, Li et al. (2024a) have crafted Chain-of-Knowledge, which dynamically pulls relevant information from a knowledge base to correct and align inconsistent rationales within CoT frameworks. These methods illustrate how strategic retrieval integration improve LLMs’ problem-solving accuracy, leading to enhanced output fidelity.
@@
Beyond the conventional types of augmentation, various domain-specific augmentations have also been applied to CoX methods for LLMs. In the realm of emotional intelligence, Lee et al. (2023b) introduce the Chain-of-Empathy, infusing psychotherapy insights to cultivate empathetic responses from LLMs. Meanwhile, Kuppa et al. (2023) propose Chain-of-Reference method that integrates legal frameworks to meticulously deconstruct and address complex legal inquiries, showcasing the versatility of CoX in specialized fields. On a similar note, Gao et al. (2024) develop a Chain-of-Abstraction framework that uses domain-specific tools to fill in abstract placeholders intentionally left in LLMs’ reasoning chains. The enhancement of linguistic tools is also evident in Chain-of-Dictionary Lu et al. (2023), which enhances machine translation with a multilingual dictionary tailored to each sentence. These diverse augmentations not only broaden the operational scope of LLMs but also underscore the potential of tailored, domain-specific enhancements.
@@ 3.3 Chain-of-Feedback
Chain-of-Feedback represents another variant of CoX. Unlike augmentation which typically precedes generation, feedback is interlaced throughout the generation process to enhance and fine-tune responses. Based on the feedback source, we categorize them as external and self-refinement feedback.
@@
Feedback from external sources provides valuable external perspectives that can guide the refinement process in LLMs. For instance, Yamada et al. (2024) introduce Chain-of-3DThought using external critiques help iteratively hone an LLM’s understanding of 3D spaces. Similarly, Wang et al. (2023b) employ a teacher-student framework in their Chain-of-Repair, where feedback from the compiler is first interpreted by a teacher LLM and then used to guide a student LLM in code generation. This approach not only corrects errors but also facilitates a learning process whereby the student model gains proficiency over time. Additionally, Liu et al. (2024a) have developed Chain-of-Hindsight, transforming direct human preference into natural language feedback that better aligns with how LLMs process information. These feedback allows for more precise refinement to model’s outputs, ensuring that responses are both accurate and contextually appropriate.
Self-Refine
The potential costs and unavailability of external feedback have led to a growing interest in self-refinement abilities within LLMs Lee et al. (2023a). Highlighted by Lei et al. (2023), Chain-of-NLI guides an LLM to evaluate and refine its outputs through a series of natural language inference tasks constructed based on its initial responses. Echoing this approach, Dhuliawala et al. (2023) introduce Chain-of-Verification empowers LLMs to self-assess through a sequence of self-generated verification questions, leading to progressively refined answers. Both methods identify and correct ungrounded outputs autonomously, enhancing the reliability of responses. Adams et al. (2023) further this concept with Chain-of-Density, which allows LLMs to iteratively incorporate self-detected missing information into their previous outputs. Together with Chain-of-SelfRevisions Le et al. (2024) and Chain-of-Feedback Ahn and Shin (2024), these frameworks exemplify how LLMs can utilize their own outputs for continuous self-improvement.
@@ 3.4 Chain-of-Models
Previous CoX methods have mostly been designed for a single LLM. Recognizing that different LLMs may have distinct specialties Xiao et al. (2024b); Xia et al. (2024b), another line of work proposes constructing a chain of models to leverage distinct strengths of each model. The Chain-of-Experts Xiao et al. (2024b) exemplifies this collaborative strategy. It involves a consortium of expert LLMs that work in sequence, each contributing its specialized knowledge to build upon the reasoning developed by its predecessors. This method is particularly effective in addressing intricate problems in operation research, where the complexity often exceeds the processing capabilities of a single LLM. Similarly, Qiu et al. (2024) deploy a chain of specialized LoRA (Low Rank Adaptation Hu et al. (2022)) networks, each fine-tuned to effectively handle different domains of a broader problem. This tailored approach ensures that specific tasks benefit from the most relevant and effective expertise, enhancing overall efficiency and outcome accuracy. In parallel, Tao et al. (2024) have developed the Chain-of-Discussion, where multiple LLMs engage in a structured dialogue, critiquing and refining each other’s contributions before reaching a consensus in the final response. This process ensures that the synthesized output is not only comprehensive but also critically evaluated from multiple perspectives.
@@ 4 Chain-of-X Tasks
As presented in the previous section, the nodes of CoX can be of various forms, enabling their applications to extend beyond LLM reasoning. This section surveys existing CoX methods categorized by tasks, as shown in Figure 2.
@@ 4.1 Multi-Modal Interaction
Although CoT was originally proposed for text generation, various CoX methods have been developed to tackle challenges in multi-modality.
@@
In the realm of vision-language models, the synergy between textual and visual data is critical Zhang et al. (2024b). CoX methods have been instrumental in enhancing this interplay. For instance, Chain-of-InstructEditing Zhang et al. (2023d) utilizes text-based instructions to guide the nuanced task of image editing, specifically for facial manipulations. This method ensures that image alterations adhere closely to textual descriptions, enhancing the accuracy and relevance of the edits. Similarly, Chain-of-Look Xi et al. (2023a) introduces a structured approach to visual entity recognition by constructing a visual semantic reasoning chain that mirrors the logical progression of CoT. This method facilitates deeper understanding and identification of visual elements through descriptive textual cues. Furthermore, Chain-of-QA Kim et al. (2024) expands this approach into a dynamic dialogue between an LLM and a visual question-answering model, tackling complex queries with a combination of textual and visual analyses. Additionally, Chain-of-Reasoning Uehara et al. (2024) and Chain-of-Manipulation Qi et al. (2024) focus on refining the process of identifying and interpreting critical details within images. These methods systematically guide the model to focus on specific regions of an image, thereby improving the model’s visual reasoning ability for more precise responses.
@@
The challenges of complex tabular data manipulation have also been studied using CoX methods. Chain-of-Command Zha et al. (2023), for example, provides LLMs with a sequence of pre-defined commands, guiding them through the accurate manipulation of tables. This structured guidance helps prevent errors that might arise from ambiguous or incorrect interpretations of the task requirements. On a related note, Chain-of-Table Wang et al. (2024) leverages tabular data as a part of the reasoning chain. Here, tables are not just data sources but act as evolving entities within the reasoning process, dynamically updating and refining themselves in response to the LLM’s queries and tasks. This iterative process allows the model to engage with the table more naturally and effectively, leading to a more nuanced understanding and manipulation of the contained information.
@@
Code generation is another task that has benefited from the introduction of CoX methods Zan et al. (2023). Chain-of-Code Li et al. (2023a), for example, tackles code generation by breaking down problems into a sequence of programs and then simulates code execution to solve the task to address the overarching task effectively. Expanding on this idea, Chain-of-Simulation La Malfa et al. (2024) takes a granular approach by executing code line by line. In contrast, Chain-of-Repair Wang et al. (2023b) draws inspiration from traditional debugging processes, where feedback from compilers is used not just to identify but also to explain bugs, facilitating a deeper learning process for the LLM as it generates fixes. Meanwhile, Chain-of-SelfRevisions Le et al. (2024) explores a creative reuse strategy, where snippets of code from previous tasks are recycled into new projects, enhancing efficiency and promoting a modular approach to code generation. Together, these methods underscore the versatility of CoX techniques in refining code generation tasks, highlighting their ability to adapt and respond to the intricacies of programming.
@@
Similarly, the field of speech generation has seen innovative applications of CoX methods. Chain-of-Information Zhang et al. (2024a), for instance, enhances speech synthesis by methodically separating and then reassembling semantic and perceptual components, which allows for more nuanced and accurate speech output. Another approach is the Chain-of-Modality Zhang et al. (2023a), which merges textual and vocal instructions to guide the speech generation process. This method not only enhances the quality of speech generation but also equips LLMs with the ability to handle conversational nuances, effectively bridging the gap between textual and speech data.
@@ 4.2 Factuality & Safety
Ensuring factual consistency and safety in LLM outputs has been critical Wang et al. (2023e); Zhang et al. (2023c); Dong et al. (2024). To make LLMs generate more factual and safer outputs, recent studies have explored the use of CoX methods in both hallucination reduction and alignment.
@@
LLMs have shown a propensity for generating hallucinations Akhtar et al. (2023); Agrawal et al. (2023); Xia et al. (2024c). Studies have explored the use of CoX methods to reduce hallucinations. For example, Chain-of-NLI Lei et al. (2023) utilizes a sequence of natural language inference problems derived from initial model outputs to guide systematic revisions, enhancing the factual accuracy of subsequent responses. Similarly, Chain-of-Verification Dhuliawala et al. (2023) prompts an LLM to generate and answer its own verification questions, enabling it to critically assess and refine its responses. Furthermore, recognizing the effectiveness of retrieval-augmented approaches in grounding responses with accurate information Gao et al. (2023), several CoX methods, e.g., Chain-of-Note Yu et al. (2023a), Chain-of-Knowledgea Li et al. (2024a), Chain-of-Actionb Pan et al. (2024), have been implemented to retrieve and integrate domain-specific knowledge at each step, effectively reducing the occurrence of incorrect or misleading information.
@@
Aligning LLMs with human preferences is another critical area where CoX methods have shown promising results Wang et al. (2023e). To enhance LLMs’ understanding of human preferences, Chain-of-Hindsight Liu et al. (2024a) transforms them into a sequence of natural language feedback for fine-tuning. Leveraging the language comprehension capabilities of LLMs, Chain-of-Hindsight achieves superior alignment performance compared to previous methods like RLHF Ouyang et al. (2022). Meanwhile, Chain-of-Utterance prompting Bhardwaj and Poria (2023) has been proposed for LLM red-teaming, establishing a jailbreaking conversation between a harmful LLM and a helpful but unsafe LLM. The harmful questions gathered through Chain-of-Utterances are utilized to create the HarmfulQA dataset, which serves as a basis for further safety alignment efforts Bhardwaj and Poria (2023). By integrating these methods, CoX frameworks not only enhance the immediate utility of LLMs but also contribute to broader efforts in developing AI systems that are both effective and ethically responsible.

### v3

@@ 3.2 Chain-of-Augmentation
While Chain-of-Intermediates method has proven effective, it falls short when LLMs have limited knowledge for specific tasks or domains. As a result, Chain-of-Augmentation has become a popular variant of CoX, where the chain is augmented with additional knowledge. Based on the types of augmented data, we categorize them as follows.
@@
Given a complex task, determining the next step can be nontrivial for LLMs even with few-shot CoT exemplars, due to misinterpretation or ambiguity Zha et al. (2023). Instructions then serve as an important augmentation, guiding LLMs through complex reasoning steps or task execution processes. For instance, Chain-of-InstructEditing Zhang et al. (2023d) harnesses this concept by generating sequential instructions to guide image editing tasks, illustrating how specific operations can refine the output for more precise editing. To avoid ambiguous user queries in table manipulation, Zha et al. (2023) introduce Chain-of-Command framework. Inferring from user instructions, it enables LLMs to employ a series of pre-defined commands for more precise table execution. In the realm of e-commerce, Li et al. (2024b) implement a similar approach Chain-of-Task, which breaks down customer interactions into manageable atomic tasks with domain-specific e-commerce instructions, significantly simplifying complex user queries.
Recently, Hayati et al. (2024) propose Chain-of-Instructions. Different from previous methods using pre-defined or human-crafted instructions, this framework iteratively uses outputs of previous steps as instructions for the next step. An instruction dataset generated is then used for fine-tuning LLMs to handle complex instructions composed of multiple subtasks. The results show that stepwise guidance can effectively improve the process and the outcomes of complex problem-solving tasks.
@@
Augmenting LLMs with historical data is essential for predictive modeling, which introduces another facet of Chain-of-Augmentation drawing contextual insights from the past. This approach is exemplified by Chain-of-Opinion Do et al. (2023), which analyzes historical user opinions to predict future reactions, offering valuable foresight into user sentiments. In user-interface exploration, Chain-of-Actiona Zhang and Zhang (2024) framework leverages past actions to guide future interactions, thereby optimizing user experience through predicted behaviors. Ma et al. (2023) take a similar approach in gaming environments like StarCraft II, where Chain-of-Summarization framework provides strategic gameplay recommendations based on a synthesis of past observations.
In addition to user modeling, the prediction of taxonomy structures also benefits from historical data, as seen in Chain-of-Layer Zeng et al. (2024), which builds upon previously identified categories. Temporal knowledge graphs receive a forward-looking treatment as well with methods like Chain-of-Historya Luo et al. (2024) and Chain-of-Historyb Xia et al. (2024e), where historical graph structures inform LLM predictions about future nodal linkages or interactions.
Retrievals
As the knowledge learned from pre-training data is limited and often outdated, LLMs frequently need to acquire external knowledge. Therefore, retrieval has become a crucial aspect of Chain-of-Augmentation. As one-step retrieval is often insufficient for complex tasks, methods have been developed to intersperse reasoning chains with explicit retrievals, thereby enhancing the quality of answers. For example, ReAct Yao et al. (2023) synergizes reasoning and acting by adaptively retrieving external knowledge to augment the reasoning chains. While ReAct prompts LLMs to make decisions, Verify-and-Edit Zhao et al. (2023) decides when to retrieve based on less-than-majority agreement consistency and corrects erroneous reasoning chains to produce a more interpretable CoT. Further refining this concept, Li et al. (2024a) develop Chain-of-Knowledgea, which dynamically pulls relevant information from both unstructured and structured knowledge sources, e.g., Wikidata and tables. While ReAct and Verify-and-Edit keep all retrieved information in the chain, Chain-of-Knowledgea makes progressive corrections and only incorporates verified retrieval results to avoid propagating misleading information.
| Chain-of-Retrievals | Self-Generated Query | Adaptive Retrieval | Retrieval Verification | Knowledge Sources |
| Self-Ask Press et al. (2023) | ✓ | Textual Corpus | ||
| ReAct Yao et al. (2023) | ✓ | ✓ | Textual Corpus | |
| Verify-and-Edit Zhao et al. (2023) | ✓ | ✓ | Textual Corpus | |
| CoKnowledgea Li et al. (2024a) | ✓ | ✓ | Textual & Tabular Data | |
| IRCoT Trivedi et al. (2023) | ✓ | Textual Corpus | ||
| CoQuery Xu et al. (2024) | ✓ | ✓ | Textual Corpus | |
| CoActionb Pan et al. (2024) | ✓ | ✓ | Textual & Market Data | |
| RAT Wang et al. (2024a) | ✓ | Textual Corpus | ||
| ToG Sun et al. (2024) | ✓ | ✓ | Knowledge Graphs | |
| GraphCoT Jin et al. (2024) | ✓ | ✓ | Knowledge Graphs |
Different from the adaptive retrieval in previous methods, another line of work explores how to compose informative queries during intermediate steps for knowledge retrieval. Press et al. (2023) propose Self-Ask, prompting LLMs to ask follow-up questions themselves and answer these sub-questions with a Google Search API before generating the final response. Similarly, IRCoT Trivedi et al. (2023), Chain-of-Question Huang et al. (2024), and RAT Wang et al. (2024a) augment each intermediate step with retrieved external knowledge iteratively refining the generations. These methods directly insert retrievals into the reasoning chains, where LLMs can only reason about a local sub-question in each generation. Thus, when there is a misleading sub-question, the entire reasoning chain afterwards will be affected. To address this limitation and ensure the coherence of the global chain, Xu et al. (2024) develop the Chain-of-Query framework, which can interactively revisit previous retrievals and make necessary reasoning direction adjustments by verifying retrieved results. Based on methodology designs and retrieval sources, we make further comparisons of previously discussed methods in Table 1.
Tools
Besides deploying a retriever to access external knowledge, recent methods have also explored utilizing other domain specific tools. MultiToolCoT Inaba et al. (2023) specifies available tools in the prompt and provides demonstration examples to enable tool using during CoT prompting. To achieve more natural tool using, ChatCoT Chen et al. (2023b) models CoT reasoning as multi-turn conversations, enabling LLMs to freely interact with tools through chatting. To further improve the efficiency of interconnected tool calls, Gao et al. (2024) develop Chain-of-Abstraction, training LLMs to decode reasoning chains with abstract placeholders to be filled with knowledge from tools. Such abstract chains enable LLMs to perform decoding and calling of external tools in parallel, thus reducing the inference delay of tool responses.
@@
Other domain-specific augmentation methods have also been explored, e.g., Chain-of-Empathy Lee et al. (2023b) for empathetic response generation, Chain-of-Reference Kuppa et al. (2023) for complex legal inquiries, and Chain-of-Dictionary Lu et al. (2023) for machine translation. These methods not only broaden the operational scope of LLMs but also underscore the potential of domain-specific CoX enhancements.
@@ 3.3 Chain-of-Feedback
Chain-of-Feedback represents another variant of CoX. Unlike augmentation which typically precedes generation, feedback is interlaced throughout the generation process to enhance and fine-tune responses. Based on the feedback source, we categorize them as external and self feedback.
@@
External feedback provides valuable perspectives overlooked by LLMs themselves. For instance, to generate 3D objects that LLMs may not have seen before, Chain-of-3DThought Yamada et al. (2024) utilizes external critiques to help iteratively hone an LLM’s understanding of 3D spaces for unconventional objects. In addition to model feedback, human feedback is another important type of external feedback especially towards aligning LLMs with human preferences. Despite the success of RLHF Ouyang et al. (2022), Chain-of-Hindsight Liu et al. (2024a) further transforms direct human preference data into natural language feedback that better aligns with how LLMs process textual information. Such feedback allows for more precise refinement to model’s outputs, ensuring that responses are both accurate and contextually appropriate.
Self Feedback
Though external feedback is critical, its costs and potential unavailability have led to a growing interest in self-refinement approaches Lee et al. (2023a). Highlighted by Madaan et al. (2023), Self-Refine first generates an initial response using an LLM and then uses the same LLM to provide feedback and refine its own response iteratively. Without any additional training, Self-Refine generates considerably better responses than direct generation. Echoing this approach, Dhuliawala et al. (2023) introduce Chain-of-Verification. Instead of directly asking LLMs to provide feedback on their own responses, Chain-of-Verification asks LLMs to first plan a series of verification questions based on the initial responses and then answer these questions themselves. After the self-assessment, LLMs are then asked to generate their final verified answers. Chain-of-NLI Lei et al. (2023) adopts similar framework but formulates a series of natural language inference problems to be answered. Similar concept has also been applied in other tasks. Chain-of-Density Adams et al. (2023) allows LLMs to iteratively refine generated summaries by incorporating self-detected missing information. Chain-of-SelfRevisions Le et al. (2024) improves modular code generation by reusing previously generated code modules.
@@ 3.4 Chain-of-Models
Previous CoX methods have mostly been designed for a single LLM. Recognizing that different LLMs may have different specialties Xiao et al. (2024b); Xia et al. (2024b), another line of work proposes constructing a chain of models to leverage distinct strengths of each model. Chain-of-Experts Xiao et al. (2024b) exemplifies this collaborative strategy. As illustrated in Figure 3, it involves a consortium of expert LLMs that work in sequence, each contributing its specialized knowledge to build upon the results developed by its predecessors. This method is particularly effective in addressing intricate problems in operations research, where the complexity often exceeds the processing capabilities of a single LLM. Similarly, Qiu et al. (2024) deploy a chain of specialized LoRA Hu et al. (2022) networks, each fine-tuned to effectively handle different domains of a broader problem. This approach ensures that specific tasks benefit from the most relevant expertise, enhancing overall efficiency and outcome accuracy. In parallel, ChatEval Chan et al. (2024) and Chain-of-Discussion Tao et al. (2024) employ multiple LLMs engaging in a dialogue, critiquing and refining each other’s contributions before reaching a consensus in the final response. This process ensures that the synthesized output is not only comprehensive but also critically evaluated from multiple perspectives.
@@ 4 Chain-of-X Tasks
CoX can be of various forms, enabling their applications in diverse areas. This section surveys CoX methods categorized by tasks as shown in Figure 2.
@@ 4.1 Multi-Modal Interaction
Real-world problem-solving tasks often involve modalities other than text. Thus, the success of CoT has drawn attention to designing frontier CoX methods for challenges in multi-modality.
@@
To handle rich features from both texts and images, the knowledge composition ability of CoX methods has been crucial in capturing key information. While MultimodalCoT Zhang et al. (2023e) incorporates image information into textual rationale generation, it lacks interaction between modalities. To address this, methods have explored using intermediate reasoning steps to infer and extract key visual information before generating final responses. For example, DDCoT Zheng et al. (2023) utilizes structured logical chains to explicitly guide the understanding of relevant image regions. Chain-of-Look Xi et al. (2023a) constructs a visual semantic reasoning chain based on textual cues for visual entity recognition. Chain-of-Manipulation Qi et al. (2024) and Chain-of-Spot Liu et al. (2024c) adopt a step-wise refinement process for identifying critical image details.
@@
Unlike texts, structured tabular data has been a challenging source for LLMs to reason with or manipulate Fang et al. (2024). To this end, CoX methods show advantages in decomposing complex table operations into a sequence of manageable subtasks. Zha et al. (2023) utilize a sequence of pre-defined commands to execute table operations step by step until the queried information is found. Taking a step further, Chain-of-Table Wang et al. (2024b) directly leverages tabular data as a part of the reasoning chain. Here, tables are not just data sources but act as evolving entities within the reasoning process, dynamically being updated in response to the LLM’s queries and tasks. This iterative process allows the model to engage with the table more naturally and effectively, leading to a more nuanced understanding and manipulation of table information.
@@
The nature of sequential execution of code generation makes it another task benefiting from CoX methods Zan et al. (2023). Chain-of-Repair Wang et al. (2023b) draws inspiration from traditional debugging processes. It employs a teacher model to interpret compiler feedback and compose a chain of code repairing steps, teaching a student model to generate debugged code. Chain-of-SelfRevisions Le et al. (2024) explores modular code generation. This method iteratively extracts and clusters sub-modules from previous generations and adds them to new reasoning chains, naturally encouraging code reuse and efficiency.
@@
The field of speech generation has also seen innovative applications of CoX methods. Chain-of-Information Zhang et al. (2024a) enhances speech synthesis by separating and then reassembling semantic and perceptual components, which allows for more nuanced and accurate speech output. Chain-of-Modality Zhang et al. (2023a) merges textual and vocal instructions to guide speech generation. This method not only enhances the quality of speech generation but also enables LLMs to handle conversational nuances, effectively bridging the gap between textual and speech data.
@@ 4.2 Factuality & Safety
Ensuring factuality and safety in LLM outputs has been critical for practical applications Wang et al. (2023g). Recent studies have also explored CoX for both hallucination reduction and alignment.
@@
LLMs have shown a propensity for generating hallucinations Agrawal et al. (2023); Xia et al. (2024c). Two main sources of hallucinations include: i) LLMs being overconfident in their incorrect understanding of the problem and overlooking details, and ii) LLMs having limited knowledge about certain tasks and generating uncertain answers Zhang et al. (2023c). For the first type, step-wise verification and iterative refinement have been applied to guide LLMs to reassess their initial response and focus on details, exemplified by Self-Refine Madaan et al. (2023), Chain-of-NLI Lei et al. (2023), and Chain-of-Verification Dhuliawala et al. (2023). For the second type, it is essential to augment LLMs with additional knowledge grounding their responses. Several CoX methods, e.g., Chain-of-Note Yu et al. (2023a), Chain-of-Knowledgea Li et al. (2024a), and Chain-of-Actionb Pan et al. (2024) retrieve domain-specific knowledge at each step, effectively reducing the occurrence of unfactual generations.
@@
Aligning LLMs with human preferences is critical to ensure that LLMs generate helpful and harmless responses. Despite the wide adoption of RLHF for LLM alignment Ouyang et al. (2022); Xia et al. (2024d), there are still challenges including the high cost of human annotation and imperfect reward functions. To help LLMs learn from any feedback form, Liu et al. (2024a) proposes transforming preference data into a sequence of natural language sentences for supervised fine-tuning. Leveraging the language comprehension capabilities of LLMs, Chain-of-Hindsight achieves better alignment performance compared to RLHF. Meanwhile, Chain-of-Utterance prompting Bhardwaj and Poria (2023) has been proposed for LLM red-teaming. It adopts a sequential structure to establish a jailbreaking conversation between a harmful LLM and a helpful but unsafe LLM, exposing safety issues of LLMs to be aligned.